A firm is must choose to buy the GSU-3300 or the UGA-3000. Both machines make the firm’s production process more efficient which in turn increases incremental cash flows. The GSU-3300 produces incremental cash flows of $25,924.00 per year for 8 years and costs $102,716.00. The UGA-3000 produces incremental cash flows of $29,487.00 per year for 9 years and cost $125,614.00. The firm’s WACC is 7.40%. What is the equivalent annual annuity of the GSU-3300?

A firm is must choose to buy the GSU-3300 or the UGA-3000. Both machines make the firm’s production process more efficient which in turn increases incremental cash flows. The GSU-3300 produces incremental cash flows of $24,975.00 per year for 8 years and costs $99,722.00. The UGA-3000 produces incremental cash flows of $29,690.00 per year for 9 years and cost $123,614.00. The firm’s WACC is 7.69%. What is the equivalent annual annuity of the UGA-3000?

Thanks in advance!
Answer #1

Case 1:

Incremental Cash flow stream for GSU-3300 as per case 1

E F G H I J 1 Case 1 WACC 7.40% 1 1 2 3 4 5 6 7 8 3 Year Incremental Cash flows for 4 GSU-3300 -1,02,716.00 25,924 25,924 25,

First we find the NPV of this cash stream. Using Excel's NPV function this is found to be

=NPV(D1,C4:J4)+B4

NPV= $49,712.78

Now using this PV, we find the PMT that we would get at the rate of WACC. This is done using the PMT function as

=PMT(D1,8,-C6)

EAA = $8,454.79 is the equivalent annual annuity of the GSU-3300

Case 2:

Incremental Cash flow stream for UGA-3000as per case 2:

C E F G H I J K AB 15 Case 2 WACC D 7.69% 16 1 2 3 4 5 6 7 8 9 17 Year Incremental Cash flows for 18 UGA-3000 -1,23,614 29,69

Similarly using the same formulas we get,

=NPV(D15,C19:K19)+B19

=$64,271

The annuity that we would get with this investment will be

=PMT(D15,9,-C21)

=$10,156.24 is the equivalent annual annuity of the UGA-3000